# Influence of capacitance and thermal fluctuations on the Josephson diode effect in asymmetric higher-harmonic SQUIDs

### v1

@@ I Introduction
While superconducting systems demonstrating nonreciprocal transport properties (the diode effect) are known for a long time [1, 2, 3], they became the focus of many studies during several last years. This superconducting diode effect (SDE) is currently actively investigated both theoretically and experimentally in various physical systems [4]. The physical mechanisms causing the SDE turn our to be quite diverse, so it can be considered as a spectacular manifestation of various fundamental physical processes. At the same time, the SDE can potentially find useful applications in superconducting electronic devices.
@@
Similar physical mechanisms [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40] can lead to the SDE in various types of Josephson junctions (JJs); in this context it is called the Josephson diode effect (JDE). This brings the rich physics of the Josephson effect [1, 2, 41] into play. Asymmetry of the Josephson effect characteristics with respect to the current direction implies realization of the JDE.
Of particular interest are SQUIDs, tunable Josephson systems of interferometer type [1, 2, 41]. A basic system of this type is shown in Fig. 1; the interferometer loop contains two JJs and is threaded by external magnetic flux $\Phi$. The up-down asymmetry of such a system (asymmetry between junctions $a$ and $b$) in the presence of magnetic flux may lead to the left-right asymmetry for the current (the JDE). This is exemplified by SQUIDs with asymmetry of effective inductances included into the two interferometer arms [42, 43, 2]. While this effect is known for a long time, miniaturization of SQUID systems diminishes inductive effects and thus suppresses this kind of the diode effect.
@@
Various SQUIDs and SQUID-like systems effectively implementing the higher-harmonic JDE mechanism have already been investigated both theoretically and experimentally [29, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. The basic quantities of interest here are the direction-dependent critical currents ($I_{c}^{+}$ and $I_{c}^{-}$) and asymmetry of the current-voltage characteristic (CVC) $I(V)$ both in the absence and in the presence of external irradiation. The CVC in Josephson systems can be described with the help of the standard resistively-shunted junction (RSJ) model [2, 41]. The JDE in SQUIDs with higher harmonics is fully captured by this model once the proper CPR is plugged into it.
A natural extension of the RSJ model is the resistively and capacitively shunted junction (RCSJ) model which takes possible capacitance (charging) effects into account [2, 41]. While the mechanical analogy of the RSJ model corresponds to strongly damped motion, the RCSJ model adds inertial effects to it. As a result, it is possible to trace crossover between overdamped and underdamped regimes. A natural question then is how the presence of capacitance influences the JDE. The RSJ model can also be extended to include a fluctuating current in order to describe thermal fluctuations [61, 62, 63], and one may expect that fluctuations lead to strong asymmetry of the CVC under the conditions of the JDE. In the context of the JDE, various charging and temperature effects have been studied before both theoretically and experimentally in Refs. [64, 65, 66, 58].
@@ II.2 RCSJ model
In order to describe the dynamics of our system, we use the RCSJ model. In this model, the Josephson equations take the following form:
@@ II.3 Asymmetric potential
To understand the origin of the JDE in our system, we use the mechanical analogy [2, 41, 67]. If we neglect the ac current and thermal fluctuations in Eq. (10), it takes the form of Newton’s equation that describes the motion of a particle with mass $\beta$ in the “washboard” potential
@@
As shown in Fig. 2, the potential shape depends on the current direction (sign of $j_{\mathrm{dc}}$) when $A\sin\tilde{\phi}\neq 0$ (as we will see later this combination is typical for asymmetries). Importantly, the value of $A$ determines the numbers of minima per period: at $A<1/4$, the potential has only one minimum per period, while at $A>1/4$, another minimum might appear at certain values of $j_{\mathrm{dc}}$.
@@ III.2 Suppression of the JDE by capacitance
In addition to numerical calculations, we investigate the suppression of the JDE by capacitance analytically. We work in the $\beta$ representation and consider the large-capacitance limit near Ohm’s law where $\overline{v}=j_{\mathrm{dc}}$ [accurate conditions of applicability will be written below, see Eqs. (32) and (33)]. We calculate the first nontrivial asymmetric correction to Ohm’s law using the “Harmonic perturbation theory” (HPT) [45]. We look for solution in the form
$\varphi(\tau)=\overline{v}\tau+\sum\limits_{n=1}^{\infty}(a_{n}\cos n\overline{v}\tau+b_{n}\sin n\overline{v}\tau).$ (29)
The phase $\varphi$ and average voltage $\overline{v}$ can be written as perturbative series
@@
where $\varphi^{(k)}$, $a^{(k)}_{n}$, and $b^{(k)}_{n}$ are contributions to the phase across junction, amplitude of the $n$th harmonic, and the average voltage in the $k$th order of the perturbation theory.
In Eq. (29), the linearly growing term is exactly equal to $\overline{v}\tau$, while the remaining part oscillates with frequencies that are multiples of $\overline{v}$.
@@ III.3 Asymmetry of the retrapping currents
In order to describe asymmetry of the retrapping currents at arbitrary $\varepsilon$ and $A$, we perform numerical calculations. The results are shown in Figs. 6 and 7. Asymmetry of the retrapping currents depends both on $A$ and $\varepsilon$, as expected. In the case of strong dissipation, $\varepsilon\gg 1$, the retrapping currents coincide with the critical currents. On the contrary, in the weak-dissipation limit, $\varepsilon\ll 1$, the retrapping currents by themselves are small, $j_{r\pm}\ll j_{c\pm}$; at the same time, asymmetry of their values is weak [see Eq. (38)]. The most interesting case is thus the regime of moderate dissipation, $\varepsilon\sim 1$, in which case the retrapping currents differ from the critical ones (at least, in one current direction) while the asymmetry is still strong.
@@ IV JDE in the presence of microwave irradiation
In this section, we discuss the CVC of the asymmetric SQUID under the influence of external microwave irradiation that generates the ac current $j_{\mathrm{ac}}\neq 0$. Except Sec. IV.3, we consider the zero-temperature limit, corresponding to $\xi=0$ in the Josephson equation in the $\beta$ representation:
@@ IV.1.2 Zero-voltage step
In the case of nonquasistationary dynamics, in order to obtain analytical results, we assume that $J(\varphi)\ll j_{\mathrm{ac}}$ and employ the perturbation theory with respect to this smallness:
@@ IV.3 Asymmetry of the resonance frequencies
As discussed in Sec. II.3, the general form of the potential and, in particular, the oscillation frequency at the well bottom $\omega_{A\pm}$, depend on the current direction. When the system is exposed to external microwave radiation, in addition to the appearance of the Shapiro steps in the R state, resonances in the S state might occur at $\tilde{\omega}\simeq\omega_{A\pm}(j_{\mathrm{dc}})$. We confine our attention to the regime of weak perturbation of the Brownian motion of the particle in the potential well at $T\neq 0$ by weak ac signal [70, 72]. In this regime, the Josephson resonance manifests itself in the great enhancement (with factors $\gamma_{\pm}$) of the thermal escape rates from the S to R state under the influence of weak ac irradiation in junctions with high quality factors $Q_{\pm}(j_{\mathrm{dc}})=\omega_{A\pm}(j_{\mathrm{dc}})\omega_{p}RC\gg 1$:
@@
$j_{\mathrm{sw}\pm}=\sqrt{1-\tilde{\omega}^{4}}+\frac{2A\tilde{\omega}^{6}\cos\tilde{\phi}}{\sqrt{1-\tilde{\omega}^{4}}}+A(1+2\tilde{\omega}^{4})\sin\tilde{\phi},$ (51)
@@ V.1 CVC at $T\neq 0$ in the zero-capacitance limit
We start our consideration from the zero-capacitance limit $\beta=0$. In this case, it is possible to obtain analytical expression for asymmetric CVC in the asymmetric potential of general form $U_{\mathrm{asym}}(\varphi)=-j_{\mathrm{dc}}\varphi-\sum\limits_{k=1}^{\infty}A_{k}\cos(k\varphi-\tilde{\phi}_{k})$. For convenience, below in this section we assume that $j_{\mathrm{dc}}>0$, and then in order to describe the negative branch of the CVC we use the symmetry
@@
In order to calculate the integrals in Eq. (58), we employ the saddle-point approximation. In the low-temperature limit, the maxima and minima of the potential are well separated from each other, and both the integrals are determined by small vicinities of the potential extrema (minima for the external integral and maxima for the internal one). We denote the locations of those maxima and minima as $\varphi_{\max\pm,m}$ and $\varphi_{\min\pm,n}$, respectively. The result of integration can then be written as
@@
In the minimal model, in the presence of only one minimum per period, one can rewrite the general expression (59) in notations of Sec. II.3 and obtain the asymmetric CVC in the following form:
@@ V.2 Escape rates at nonzero capacitance
In order to emphasize the physical meaning of $\tilde{\tau}_{l\pm}^{-1}$, we note that it is nothing but the escape rates $\Gamma_{\pm}$ entering Eq. (50). These rates are also related to the average voltage by the simple expression in the overdamped regime:
@@ V.3 Asymmetry of the switching currents
In order to obtain explicit expressions for the switching currents, one should substitute here asymptotic expressions for $\Delta U_{\pm}$ and $\tilde{\omega}_{\mathrm{att}\pm}$ and then solve the resulting transcendental equation.
@@ VI Discussion
As discussed in Sec. IV, when the system is exposed to small external ac current, the Josephson plasma resonances in junctions with $\varepsilon\ll 1$ arise when the external frequency satisfies the resonance condition $\tilde{\omega}\simeq\omega_{A\pm}(j_{\mathrm{dc}})$. In this case, the ac current can greatly enhance the thermal escape rate for one current direction (which satisfies the resonance condition) and thus stimulate switching from the S to R state, while leaving the system in the S state for the opposite current direction (which does not satisfy the resonance condition). As shown in Fig. 3, this behavior occurs at $|j_{\mathrm{dc}}|<j_{c\pm}$. Therefore, this makes it possible to expand the operational range of the Josephson diodes when external ac current is applied. To this end, it is preferable to apply frequency $\tilde{\omega}$ in resonance with the largest of the two values $\omega_{A\pm}$. This is due to the fact that the enhancement factor of thermal escapes $\gamma(\tilde{\omega})$ rapidly decreases at $\tilde{\omega}>\omega_{A}(j_{\mathrm{dc}})$ [72, 70]. The above procedure helps avoid parasitic switching from the S to R state in the opposite current direction.
@@ A.1 CVC in the large-capacitance limit
In order to apply the HPT in the large-capacitance limit, we represent the phase and voltage as series, see Eqs. (29) and (30). Then, we substitute these expansions to Eq. (23), expand the equation into the Fourier series, and solve it in the required order of the perturbation theory assuming conditions (32).
@@ A.2 First Shapiro steps
In order to demonstrate how the HPT works in this case, we calculate the heights of the first Shapiro steps at $\beta=0$. We emphasize that at $\beta\neq 0$, the leading asymmetric term in the heights will be the same. Similarly, the results for the large-capacitance limit can be obtained by this technique.
@@ A.2.1 HPT for the first Shapiro steps at $\beta=0$
$\overline{v}(-a^{(1)}_{1}\sin\overline{v}\tau+a^{(1)}_{1}\cos\overline{v}\tau+a^{(1)}_{2}\cos 2\overline{v}\tau+b^{(1)}_{2}\sin 2\overline{v}\tau)=j^{(1)}+j_{\mathrm{ac}}\cos(\overline{v}\tau+\delta)-\sin\overline{v}\tau-A\sin(2\overline{v}\tau-\tilde{\phi}),$ (84)
@@
$\varphi^{(1)}=\left[j_{\mathrm{ac}}\sin(\overline{v}\tau+\delta)+\cos\overline{v}\tau+(A/2)\cos(2\overline{v}\tau-\tilde{\phi})\right]/\overline{v}.$ (85)
@@
As a result, in the leading order of the HPT, asymmetry of the heights of the first Shapiro steps takes form (48). Note that in this equation, $j_{\mathrm{ac}}$ and $A$ are assumed to be small only compared to large $\omega$.
@@ B.2 First order with respect to $\varepsilon$ and to $A$
Equation (96) takes into account the corrections to the retrapping current in the first order with respect to $A$. However, the retrapping current in this order is symmetric and, as one can check, this will be so in any order with respect to $A$ in the first order with respect to $\varepsilon$ [because in the first order of the perturbation theory with respect to $\varepsilon$ we neglect the dc-current contribution in Eq. (35)]. Therefore, we need to consider the next order of the perturbation theory in order to find asymmetry of the retrapping currents $j_{r+}\neq j_{r-}$.

### v2

@@ I Introduction
While superconducting systems demonstrating nonreciprocal transport properties (the diode effect) have been known for a long time [1, 2, 3], they have become the focus of many studies during recent years. This superconducting diode effect (SDE) is currently actively investigated both theoretically and experimentally in various physical systems [4]. The physical mechanisms causing the SDE turn out to be quite diverse, so it can be considered as a spectacular manifestation of various fundamental physical processes. At the same time, the SDE can potentially find useful applications in superconducting electronic devices.
@@
Similar physical mechanisms [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] can lead to the SDE in various types of Josephson junctions (JJs); in this context it is called the Josephson diode effect (JDE). This brings the rich physics of the Josephson effect [1, 2, 44] into play. Asymmetry of the Josephson effect characteristics with respect to the current direction implies realization of the JDE.
Of particular interest are SQUIDs, tunable Josephson systems of interferometer type [1, 2, 44]. A basic system of this type is shown in Fig. 1; the interferometer loop contains two JJs and is threaded by external magnetic flux $\Phi$. The up-down asymmetry of such a system (asymmetry between junctions $a$ and $b$) in the presence of magnetic flux may lead to the left-right asymmetry for the current (the JDE). This is exemplified by SQUIDs with asymmetry of effective inductances included into the two interferometer arms [45, 46, 2]. While this effect has been known for a long time, miniaturization of SQUID systems diminishes inductive effects and thus suppresses this kind of the diode effect.
@@
Various SQUIDs and SQUID-like systems effectively implementing the higher-harmonic JDE mechanism have already been investigated both theoretically and experimentally [29, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. The basic quantities of interest here are the direction-dependent critical currents ($I_{c}^{+}$ and $I_{c}^{-}$) and asymmetry of the current-voltage characteristic (CVC) $I(V)$ both in the absence and in the presence of external irradiation. The CVC in Josephson systems can be described with the help of the standard resistively shunted junction (RSJ) model [2, 44]. The JDE in SQUIDs with higher harmonics is fully captured by this model once the proper CPR is plugged into it.
A natural extension of the RSJ model is the resistively and capacitively shunted junction (RCSJ) model which takes possible capacitance (charging) effects into account [2, 44]. While the mechanical analogy of the RSJ model corresponds to strongly damped motion, the RCSJ model adds inertial effects to it. As a result, it is possible to trace crossover between overdamped and underdamped regimes. A natural question then is how the presence of capacitance influences the JDE. The RSJ model can also be extended to include a fluctuating current to describe thermal fluctuations [64, 65, 66], and one may expect that fluctuations lead to strong asymmetry of the CVC under the conditions of the JDE. In the context of the JDE, various charging and temperature effects have been studied before both theoretically and experimentally in Refs. [67, 68, 69, 61].
@@ II.2 RCSJ model
To describe the dynamics of our system, we use the RCSJ model. In this model, the Josephson equations take the following form:
@@ II.3 Asymmetric potential
To understand the origin of the JDE in our system, we use the mechanical analogy [2, 44, 70]. If we neglect the ac current and thermal fluctuations in Eq. (10), then it takes the form of Newton’s equation that describes the motion of a particle with mass $\beta$ in the “washboard” potential,
@@
As shown in Fig. 2, the potential shape depends on the current direction (sign of $j_{\mathrm{dc}}$) when $A\sin\tilde{\phi}\neq 0$ (as we will see later this combination is typical for asymmetries). Importantly, the value of $A$ determines the numbers of minima per period: at $A<1/4$, the potential has only one minimum per period, while at $A>1/4$, another minimum appears at certain values of $j_{\mathrm{dc}}$.
@@ III.2 Suppression of the JDE by capacitance
In addition to numerical calculations, we investigate the suppression of the JDE by capacitance analytically. We work in the $\beta$ representation and consider the large-capacitance limit near Ohm’s law where $\overline{v}=j_{\mathrm{dc}}$ [accurate conditions of applicability will be written below, see Eqs. (32) and (33)]. We calculate the first nontrivial asymmetric correction to Ohm’s law using the “harmonic perturbation theory” (HPT) [48].
Generally, the solution can be represented in the form
$\varphi(\tau)=\overline{v}\tau+\sum\limits_{n=1}^{\infty}(a_{n}\cos n\overline{v}\tau+b_{n}\sin n\overline{v}\tau),$ (29)
where we take into account that the slope of the linearly growing term is exactly equal to $\overline{v}$, while the remaining part oscillates with frequencies that are multiples of $\overline{v}$ (which is a manifestation of the voltage periodicity with period $2\pi/\overline{v}$).
While representation (29) is general, its parameters $a_{n}$, $b_{n}$, and $\overline{v}$ have to be determined from a complicated system of equations, which can be done only numerically. At the same time, in certain limiting cases, they can be found perturbatively:
@@
where $\varphi^{(k)}$, $a^{(k)}_{n}$, $b^{(k)}_{n}$, and $\overline{v}^{(k)}$ are contributions to the phase across junction, amplitudes of the $n$th harmonics, and the average voltage in the $k$th order of the perturbation theory. The HPT that we employ below is thus a way to find perturbative expansions (30) for the parameters of the general harmonic representation (29).
@@ III.3 Asymmetry of the retrapping currents
To describe asymmetry of the retrapping currents at arbitrary $\varepsilon$ and $A$, we perform numerical calculations. The results are shown in Figs. 6 and 7. Asymmetry of the retrapping currents depends both on $A$ and $\varepsilon$, as expected. In the case of strong dissipation, $\varepsilon\gg 1$, the retrapping currents coincide with the critical currents. On the contrary, in the weak-dissipation limit, $\varepsilon\ll 1$, the retrapping currents by themselves are small, $j_{r\pm}\ll j_{c\pm}$; at the same time, asymmetry of their values is weak [see Eq. (38)]. The most interesting case is thus the regime of moderate dissipation, $\varepsilon\sim 1$, in which case the retrapping currents differ from the critical ones (at least, in one current direction) while the asymmetry is still strong.
@@ IV JDE in the presence of microwave irradiation
In this section, we discuss the CVC of the asymmetric SQUID under the influence of external microwave irradiation that generates the ac current $j_{\mathrm{ac}}\neq 0$. Except in Sec. IV.3, we consider the zero-temperature limit, corresponding to $\xi=0$ in the Josephson equation in the $\beta$ representation:
@@ IV.1.2 Zero-voltage step
In the case of nonquasistationary dynamics, to obtain analytical results, we assume that $J(\varphi)\ll j_{\mathrm{ac}}$ and employ the perturbation theory with respect to this smallness:
@@ IV.3 Asymmetry of the resonance frequencies
As discussed in Sec. II.3, the general form of the potential and, in particular, the oscillation frequency at the well bottom $\omega_{A\pm}$, depend on the current direction. When the system is exposed to external microwave radiation, in addition to the appearance of the Shapiro steps in the R state, resonances in the S state might occur at $\tilde{\omega}\simeq\omega_{A\pm}(j_{\mathrm{dc}})$. We confine our attention to the regime of weak perturbation of the Brownian motion of the particle in the potential well at $T\neq 0$ by weak ac signal [75, 76, 73]. In this regime, the Josephson resonance manifests itself in the great enhancement (with factors $\gamma_{\pm}$) of the thermal escape rates from the S to R state under the influence of weak ac irradiation in junctions with high quality factors $Q_{\pm}(j_{\mathrm{dc}})=\omega_{A\pm}(j_{\mathrm{dc}})\omega_{p}RC\gg 1$:
@@
$j_{\mathrm{sw}\pm}=\sqrt{1-\tilde{\omega}^{4}}+\frac{2A\tilde{\omega}^{6}\cos\tilde{\phi}}{\sqrt{1-\tilde{\omega}^{4}}}\pm A(1+2\tilde{\omega}^{4})\sin\tilde{\phi},$ (51)
@@ V.1 CVC at $T\neq 0$ in the zero-capacitance limit
We start our consideration from the zero-capacitance limit $\beta=0$. In this case, it is possible to obtain analytical expression for asymmetric CVC in the asymmetric potential of general form $U_{\mathrm{asym}}(\varphi)=-j_{\mathrm{dc}}\varphi-\sum\limits_{k=1}^{\infty}A_{k}\cos(k\varphi-\tilde{\phi}_{k})$. For convenience, below in this section we assume that $j_{\mathrm{dc}}>0$, and then to describe the negative branch of the CVC we use the symmetry
@@
To calculate the integrals in Eq. (58), we employ the saddle-point approximation. In the low-temperature limit, the maxima and minima of the potential are well separated from each other, and both the integrals are determined by small vicinities of the potential extrema (minima for the external integral and maxima for the internal one). We denote the locations of those maxima and minima as $\varphi_{\max\pm,m}$ and $\varphi_{\min\pm,n}$, respectively. The result of integration can then be written as
@@
In the minimal model, in the presence of only one minimum per period, [^1] In the presence of two nontrivial minima per period, the stochastic dynamics become more complicated due to interplay between different S states of the junction [78]. one can rewrite the general expression (59) in notations of Sec. II.3 and obtain the asymmetric CVC in the following form:
@@ V.2 Escape rates at nonzero capacitance
To emphasize the physical meaning of $\tilde{\tau}_{l\pm}^{-1}$, we note that it is nothing but the escape rates $\Gamma_{\pm}$ entering Eq. (50). These rates are also related to the average voltage by the simple expression in the overdamped regime:
@@ V.3 Asymmetry of the switching currents
To obtain explicit expressions for the switching currents, one should substitute here asymptotic expressions for $\Delta U_{\pm}$ and $\tilde{\omega}_{\mathrm{att}\pm}$ and then solve the resulting transcendental equation.
@@ VI Discussion
As discussed in Sec. IV, when the system is exposed to small external ac current, the Josephson plasma resonances in junctions with $\varepsilon\ll 1$ arise when the external frequency satisfies the resonance condition $\tilde{\omega}\simeq\omega_{A\pm}(j_{\mathrm{dc}})$. In this case, the ac current can greatly enhance the thermal escape rate for one current direction (which satisfies the resonance condition) and thus stimulate switching from the S to R state, while leaving the system in the S state for the opposite current direction (which does not satisfy the resonance condition). As shown in Fig. 3, this behavior occurs at $|j_{\mathrm{dc}}|<j_{c\pm}$. Therefore, this makes it possible to expand the operational range of the Josephson diodes when external ac current is applied. To this end, it is preferable to apply frequency $\tilde{\omega}$ in resonance with the largest of the two values $\omega_{A\pm}$. This is due to the fact that the enhancement factor of thermal escapes $\gamma(\tilde{\omega})$ rapidly decreases at $\tilde{\omega}>\omega_{A}(j_{\mathrm{dc}})$ [75, 76, 73]. The above procedure helps avoid parasitic switching from the S to R state in the opposite current direction.
@@ A.1 CVC in the large-capacitance limit
To apply the HPT in the large-capacitance limit, we represent the phase and voltage as series, see Eqs. (29) and (30). Then, we substitute these expansions to Eq. (23), expand the equation into the Fourier series, and solve it in the required order of the perturbation theory assuming conditions (32).
@@ A.2 First Shapiro steps
To demonstrate how the HPT works in this case, we calculate the heights of the first Shapiro steps at $\beta=0$. We emphasize that at $\beta\neq 0$, the leading asymmetric term in the heights will be the same. Similarly, the results for the large-capacitance limit can be obtained by this technique.
@@ A.2.1 HPT for the first Shapiro steps at $\beta=0$
$\overline{v}(-a^{(1)}_{1}\sin\overline{v}\tau+b^{(1)}_{1}\cos\overline{v}\tau-2a^{(1)}_{2}\sin 2\overline{v}\tau+2b^{(1)}_{2}\cos 2\overline{v}\tau)=j^{(1)}+j_{\mathrm{ac}}\cos(\overline{v}\tau+\delta)-\sin\overline{v}\tau-A\sin(2\overline{v}\tau-\tilde{\phi}),$ (84)
@@
$\varphi^{(1)}=\left[j_{\mathrm{ac}}\sin(\overline{v}\tau+\delta)+\cos\overline{v}\tau+(A/2)\cos(2\overline{v}\tau-\tilde{\phi})\right]/\overline{v},\qquad j^{(1)}=0.$ (85)
@@
As a result, in the leading order of the HPT, asymmetry of the heights of the first Shapiro steps takes form of Eq. (48). Note that in this equation, $j_{\mathrm{ac}}$ and $A$ are assumed to be small only compared to large $\omega$.
@@ B.2 First order with respect to $\varepsilon$ and to $A$
Equation (96) takes into account the corrections to the retrapping current in the first order with respect to $A$. However, the retrapping current in this order is symmetric and, as one can check, this will be so in any order with respect to $A$ in the first order with respect to $\varepsilon$ [because in the first order of the perturbation theory with respect to $\varepsilon$ we neglect the dc-current contribution in Eq. (35)]. Therefore, we need to consider the next order of the perturbation theory to find asymmetry of the retrapping currents $j_{r+}\neq j_{r-}$.